OpenStudy (anonymous):

PLEASE HELP w/ THIS WORD PROBLEM: A telephone company offers two long-distance plans: Plan A:$25 per month and 5¢ per minute Plan B:$5 per month and 12¢ per minute For how many minutes of long-distance calls would Plan B be financially advantageous?

OpenStudy (anonymous):

let x = minutes you spend so the equations here: 5+0.12x = 25 + 0.05x where the total amount in plan B has to be less than plan A solve for x 20 = 0.07x x = 20/0.07 x = 285.7 round to 285 so if you use 285 or less minutes, plan B is better

OpenStudy (anonymous):

Thank you! this helped a lot